OpenStudy (anonymous):

Write an appropriate direct variation equation if y = 8 when x = 2.

OpenStudy (anonymous):

A. y = -3x + 1 B. y = -3x C. y = 3x D. y = 3x + 1

OpenStudy (anonymous):

none

OpenStudy (anonymous):

ha well that's not an anser lol

OpenStudy (anonymous):

oh oh hold on wrong one's

OpenStudy (anonymous):

well, put x=2 and see none of the options gives you 8

OpenStudy (anonymous):

B. y = 4x C. y = 8x

OpenStudy (anonymous):

B lol

OpenStudy (anonymous):

hope that helps ya , genius soul

OpenStudy (anonymous):

\[\frac{ }{ x}\] \[y=\frac{ 16 }{ x}\]

OpenStudy (anonymous):

ugh the top one is 4/x GEEz!

OpenStudy (anonymous):

again y=16/x gives you 8 when x=2

OpenStudy (anonymous):

so it's that one?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

16/x? lol b/c you said 4x the first time lol

OpenStudy (anonymous):

in both the cases if your gona put x=2 you'll get y=8

OpenStudy (anonymous):

oh that makes sense so it is that one16/x or it can be both?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

I can't choose both lol

OpenStudy (anonymous):

then go for y=4x, thats gona give you linear relationship

OpenStudy (anonymous):

okie dokie

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

I have so many more..